OpenStudy (anonymous):

Help solving 2x3xnxn=54

ganeshie8 (ganeshie8):

start by dividing `2` both sides

ganeshie8 (ganeshie8):

\[\large \begin{align} 2\times 3\times n\times n&=54\\~\\ \dfrac{2\times 3\times n\times n}{2}&=\dfrac{54}{2}\end{align}\]